# Temporal Interest Network for User Response Prediction

### v3

@@ Abstract.
User behaviors are among the most critical features for user response prediction in recommendation systems. Many works have revealed that a user’s behavior reflects her interest in the candidate item, owing to the semantic or temporal correlation between them. While the literature has individually examined each of these correlations, researchers have yet to analyze them in combination, that is, the semantic-temporal correlation. We empirically measure this correlation and observe intuitive yet robust patterns. We then examine several popular user interest models and find that, surprisingly, none of them learn such correlation well.
To fill this gap, we propose a Temporal Interest Network (TIN) to capture the semantic-temporal correlation simultaneously between behaviors and the target. We achieve this by incorporating target-aware temporal encoding, in addition to semantic encoding, to represent behaviors and the target. Furthermore, we conduct explicit 4-way interaction by deploying target-aware attention and target-aware representation to capture both semantic and temporal correlation. We conduct comprehensive evaluations on two popular public datasets, and our proposed TIN outperforms the best-performing baselines by 0.43% and 0.29% on GAUC, respectively. During online A/B testing in Tencent’s advertising platform, TIN achieves 1.65% cost lift and 1.93% GMV lift over the base model. It has been successfully deployed in production since October 2023, serving the WeChat Moments traffic. We have released our code at https://github.com/zhouxy1003/TIN.
@@
journalyear: 2024copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589335.3648340isbn: 979-8-4007-0172-6/24/05ccs: Computing methodologies Semantic networks
@@ 1. Introduction
In recent decades, users have been confronted with an overwhelming amount of information while browsing websites or mobile apps. This poses significant challenges for electronic retailers, content providers, and online advertising platforms in their quest to recommend suitable items to individual users within specific contexts. Consequently, recommendation systems have gained widespread deployment to capture users’ interests and predict their preferences from an extensive pool of candidate items. For instance, in cost-per-click (CPC) advertising, advertising platforms must bid for each ad based on the estimated value of the impression, which relies on the bid value and the estimated Click-Through Rate (CTR). As a result, the accurate prediction of user response becomes a critical factor and has garnered considerable research attention.
@@
Figure 1. (a): The ground truth Semantic-Temporal Correlation (STC) between history behaviors of Top-5 categories (y-axis) at various target-relative positions from 1 to 10 (x-axis), and the target category of 674; (b,c,d): the learned STC in DIN, SASRec and BST. Numbers in parentheses denote the Pearson correlation coefficient between ground truth and learned correlation on behaviors with the same category as the target (3rd row).
@@
Intuitively, a user’s interest is reflected by what items she has interacted with and when these interactions happened. For example, if she has clicked ads on video games many times, then it’s a good idea to recommend ads for new video games to her in the future. However, if all her clicks on video games happened 1 year ago, but she clicked on clothing ads several times recently, then it’s better to recommend clothing ads. Therefore, when capturing a user’s interest in a candidate item, it’s crucial to consider both the semantic correlation and the temporal correlation between behaviors and the candidate. To our surprise, there is neither a study measuring such semantic-temporal correlation nor an examination of how well existing methods capture it.
Motivated by the lack of such analysis, we propose an empirical measurement of the ground truth semantic-temporal correlation by mutual information. On the Amazon dataset, which contains users’ reviews on items, we calculate the mutual information between history reviews with Top-5 categories at various target-relative positions and the target item with a specific category. As illustrated in Fig. 1(a), our results reveal a compelling semantic-temporal pattern. We then examined the learned correlation of several popular user interest models such as DIN (Zhou et al., 2018), SASRec (Kang and McAuley, 2018) and BST (Chen et al., 2019), and show the results in Fig. 1. To our surprise, none of these models is capable of capturing the semantic-temporal pattern well.
@@
Each of the three components in TIN is critical in acquiring the semantic-temporal correlation. To empirically validate the significance of each component, we conduct an ablation study on the Amazon dataset. The ablation of each component results in a notable decrease in the GAUC by 0.99%, 0.73%, and 0.61%, respectively. It is noteworthy that all existing methods can be considered as ablated variants of TIN, as they lack one or more of these critical components. The contribution of this paper can be summarized as follows:
@@ 2. Semantic-Temporal Correlation
This section presents a novel measurement to quantize the ground truth semantic-temporal correlation between behaviors and the target via mutual information. We then examine several popular user interest models on their ability to learn such correlation.
@@ 2.2. How do Existing Methods Capture Semantic-Temporal Pattern?
We observe notable disparities in the learned correlations of these models compared to the ground truth one. DIN fails to capture any temporal pattern, as evidenced by the absence of decay along positions. SASRec and BST demonstrate limited proficiency in learning temporal correlation, as there is a slight decay along positions. However, they fall short in capturing the semantic correlation, as they learn high correlations for many other categories besides target category 674 (3rd row). Existing methods’ failure to learn semantic-temporal correlation motivates us to develop a more effective user interest model. For a more comprehensive understanding of how we measure the learned correlation of each model, please refer to Sec. 4.3.
@@ 3. Temporal Interest Network
Figure 2. The left part illustrates the whole architecture of TIN; the right part illustrates the architecture of the Temporal Interest Module, which introduces target-aware temporal encoding to preserve the temporal information, and target-aware attention and representation for explicit 4-way interaction.
@@ 3.1. Temporal Interest Module
(2) $\bm{u}_{\text{TIN}}=\sum_{i\in\mathcal{H}}\alpha(\tilde{\bm{e}}_{i},\tilde{\bm{v}}_{t})\cdot(\tilde{\bm{e}}_{i}\odot\tilde{\bm{v}}_{t}),$
@@ 3.1.2. Target-aware Attention and Target-aware Representation
Building upon target attention methods (Zhou et al., 2018, 2019; Feng et al., 2019), we aim to capture the importance of each behavior $X_{i}$ through an attention function. In this regard, we adopt the Scaled Dot-Product (Vaswani et al., 2017; Zheng et al., 2022) as our activation function: $\alpha(\tilde{\bm{e}}_{i},\tilde{\bm{v}}_{t})=\sigma(\frac{\langle\tilde{\bm{v}}_{t},\tilde{\bm{e}}_{i}\rangle}{\sqrt{d}})$, where $\sigma(z)$ denotes the softmax function over all behaviors of a given user.
@@ 3.3. Connections to Transformer
They differ from each other regarding: 1. Scaled Dot-Product Attention is applied between each pair of tokens within the sequence, while TIM is only applied between the target and each item in the sequence, that is, Scaled Dot-Product Attention is a Self-Attention upon a sequence, while TIM is a Target-Attention upon a sequence and a target; 2. The representation in Scaled Dot-Product Attention is 1st-order only, that is, $V$, while it’s 2nd-order in TIM, that is, $(\tilde{\bm{e}}_{i}\odot\tilde{\bm{v}}_{t})$.
@@ 4. Experiments
RQ1: How does the performance of our proposed Temporal Interest Network (TIN) compare to state-of-the-art methods?
@@
RQ4: How do various temporal encoding methods perform in comparison to each other?
@@ 4.1. RQ1: Overall Performance Evaluation
The performance evaluation results are presented in Table 1. It is evident that Avg Pooling & Concat exhibits the poorest performance on both datasets. This can be attributed to its lack of all three essential components, rendering it incapable of capturing either temporal correlation or semantic correlation (feature interaction). In contrast, Avg Pooling & Product demonstrates a significant performance improvement compared to Sum Pooling & Concat, with gains of 9.2e-3 and 5.3e-3 on each dataset, respectively. This improvement can be attributed to the inclusion of target-aware representation through the product operation, as discussed in (Rendle et al., 2020). While DIN’ performs worse than Avg Pooling & Product on three metrics, except for GAUC on the Alibaba dataset, this can be attributed to the limited expressiveness of Temporal Attention (TA) compared to TR. DIN outperforms DIN’ due to its incorporation of target-aware representation.
@@ 4.3. RQ3: Measurement of Learned Correlation
Figure 3. Learned semantic-temporal correlation on the Amazon dataset. The numbers in parentheses represent the Pearson correlation coefficient between the ground truth (as shown in Fig. 1(a)) and the learned correlation on behaviors with the same category as the target (3rd row).
@@ 4.4. RQ4: Temporal Encoding Methods
To evaluate these two temporal encoding methods, we partitioned the Amazon dataset based on the sequence length and conducted a comparative analysis. As shown in Table 3, TTE-P consistently outperforms COE across all sequence length ranges, with the performance gap widening for longer sequences. Unfortunately, we are unable to perform a similar comparison on the Alibaba dataset since most sequences are padded to a fixed length of 50.
Next, we present the comparison results of TTE-P, TTE-T, and Timestamp Encoding on both datasets in Table 4. Specifically, TTE-T split the time interval into 10 buckets with equal frequency binning function. Upon replacing TTE-P with TTE-T, we observe an increase in the GAUC by 8.6e-3 on the Amazon dataset and a slight drop of 0.8e-3 on the Alibaba dataset. This discrepancy can be attributed to the distinct time ranges of behaviors in the two datasets. The Amazon dataset encompasses review data from 1996 to 2014, whereas the Alibaba dataset comprises user behaviors recorded over a continuous period of 8 days. The larger time range of the Amazon dataset renders TTE-T more informative than TTE-P. Conversely, on the Alibaba dataset with a relatively short time range, TTE-P proves to be more effective in capturing temporal patterns.
@@ 4.6. Online A/B Testing
We apply TIN to the user’s 2-year clicked ads category sequence feature in Tencent’s WeChat Moments, and use both TTE-P and TTE-T to capture the temporal correlation. The backbone model employs a Heterogeneous Experts with Multi-Embedding architecture (Guo et al., 2023; Su et al., 2023b; Pan et al., 2024). Specifically, we learn multiple different feature interaction experts, e.g., GwPFM (Pan et al., 2024) (a variant of FFM (Juan et al., 2017) and FwFM (Pan et al., 2018)), IPNN (Qu et al., 2016), DCN V2 (Wang et al., 2021), or FlatDNN. There are hundreds of features in total, mostly user behavior features, along with ads and context side features. Multiple embedding tables are learned for all features, each corresponding to one or several experts. TIN functions as an additional expert alongside the existing three experts, sharing embeddings with GwPFM and FlatDNN.
We conduct online A/B testing from September 2023 to October 2023. In the baseline, the user interest is modeled by feeding both user behavior features as well as target ads features to GwPFM or IPNN. We previously attempted to use DIN, but it did not surpass the baseline’s performance. This is possibly due to that explicit interaction models such as GwPFM or IPNN are already able to learn the semantic correlation between behaviors and the target. During the two-week 20% A/B testing, TIN demonstrated promising results, achieving a 1.65% increase in cost and a 1.93% increase in GMV (Gross Merchandise Value) compared to the baseline. These improvements were statistically significant according to t-tests. Based on our estimation, these performance gains could potentially lead to a revenue increase of hundreds of millions of dollars per year. TIN has been successfully deployed on production in several scenarios.
@@ 6. Conclusion
7. Acknowledgement
We gratefully acknowledge the contributions of the following: Xueming Qiu, Lei Mu, Xian Hu, Yaqian Zhang, Ming Yue, Wenbo Liu, Xiaobo Li, Zezhen Xu, Jiayu Sun, Xun Liu, Yi Li, Ximei Wang, Junwen Cheng, Yan Tan, Yuxiong Li, Zhaohua Li, Kuo Zhang and Yufei Zheng.
This work was supported by the National Natural Science Foundation of China [U23A20309, 62272302, 62172276, 62372296], Shanghai Municipal Science and Technology Major Project [2021SHZDZX-0102], and the Tencent Rhinoceros Project. Haolin Zhou, Xinyi Zhou, Xiaofeng Gao and Guihai Chen are in the MoE Key Lab of Artificial Intelligence, Department of Computer Science and Engineering, Shanghai Jiao Tong University, Shanghai, China. Xiaofeng Gao is the corresponding author.
@@ A.2. Expressiveness Study of Target-aware Attention and Target-aware Representation
Analysis of semantic-temporal correlation on more categories

### v4

@@ Abstract.
User behaviors are among the most critical features for user response prediction in recommendation systems. Many works have revealed that a user’s behavior reflects her interest in the candidate item, owing to their semantic or temporal correlation. While the literature has individually examined each of these correlations, researchers have yet to analyze them in combination, that is, the semantic-temporal correlation. We empirically measure this correlation and observe intuitive yet robust patterns. We then examine several popular user interest models and find that, surprisingly, none of them learn such correlation well.
To fill this gap, we propose a Temporal Interest Network (TIN) to capture the semantic-temporal correlation simultaneously between behaviors and the target. We achieve this by incorporating target-aware temporal encoding, in addition to semantic encoding, to represent behaviors and the target. Furthermore, we conduct explicit 4-way interaction by deploying target-aware attention and target-aware representation to capture both semantic and temporal correlation. We conduct comprehensive evaluations on two popular public datasets, and our proposed TIN outperforms the best-performing baselines by 0.43% and 0.29% on GAUC, respectively. During online A/B testing in Tencent’s advertising platform, TIN achieves 1.65% cost lift, and 1.93% GMV lift over the base model. It has been successfully deployed in production since October 2023, serving the WeChat Moments traffic. We have released our code at https://github.com/zhouxy1003/TIN.
@@
journalyear: 2024copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589335.3648340isbn: 979-8-4007-0172-6/24/05ccs: Information systems Display advertisingccs: Computing methodologies Neural networksccs: Computing methodologies Factorization methods
@@ 1. Introduction
In recent decades, users have been confronted with overwhelming information while browsing websites or mobile apps. This poses significant challenges for electronic retailers, content providers, and online advertising platforms in their quest to recommend suitable items to individual users within specific contexts. Consequently, recommendation systems have gained widespread deployment to capture users’ interests and predict their preferences from an extensive pool of candidate items. For instance, in cost-per-click (CPC) advertising, advertising platforms must bid for each ad based on the estimated value of the impression, which relies on the bid value and the estimated Click-Through Rate (CTR). As a result, the accurate prediction of user response becomes a critical factor and has garnered considerable research attention.
@@
Figure 1. (a): The ground truth Semantic-Temporal Correlation (STC) between history behaviors of Top-5 frequent categories (y-axis) at various target-relative positions from 1 to 10 (x-axis), and the target category of 674; (b,c,d): the learned STC in DIN, SASRec and BST. Numbers in parentheses denote the Pearson correlation coefficient between ground truth STC and learned STC on behaviors with the same category as the target (3rd row in each figure), i.e., 674.
@@
Intuitively, a user’s interest is reflected by what items she has interacted with and when these interactions happened. For example, if she has clicked ads on video games many times, then it’s a good idea to recommend ads for new video games to her in the future. However, if all her clicks on video games happened 1 year ago, but she clicked on clothing ads several times recently, then it’s better to recommend clothing ads. Therefore, when capturing a user’s interest in a candidate item, it’s crucial to consider both the semantic and temporal correlations between behaviors and the candidate. To our surprise, no study measures such semantic-temporal correlation or examines how well existing methods capture it.
Motivated by the lack of such analysis, we propose an empirical measurement of the ground truth semantic-temporal correlation by mutual information. On the Amazon dataset, which contains users’ reviews on items, we calculate the mutual information between history reviews with Top-5 categories at various target-relative positions and the target item with a specific category. As illustrated in Fig. 1(a), our results reveal a compelling semantic-temporal pattern. We then examined the learned correlation of several popular user interest models such as DIN (Zhou et al., 2018), SASRec (Kang and McAuley, 2018) and BST (Chen et al., 2019), and show the results in Fig. 1. To our surprise, none of these models can capture the semantic-temporal pattern well.
@@
Each of the three components in TIN is critical in acquiring the semantic-temporal correlation. To empirically validate the significance of each component, we conduct an ablation study on the Amazon dataset. The ablation of each component results in a notable decrease in the GAUC by 0.99%, 0.73%, and 0.61%, respectively. Notably, all existing methods can be considered ablated variants of TIN, as they lack one or more of these critical components. The contribution of this paper can be summarized as follows:
@@ 2. Semantic-Temporal Correlation
This section presents a novel measurement to quantize the ground truth semantic-temporal correlation between behaviors and the target via mutual information. We then examine several popular user interest models’ ability to learn such correlation.
@@ 2.2. How do Existing Methods Capture Semantic-Temporal Pattern?
We observe notable disparities in the learned correlations of these models compared to the ground truth one. DIN fails to capture any temporal pattern, as evidenced by the absence of decay along positions. SASRec and BST demonstrate limited proficiency in learning temporal correlation, as there is a slight decay along positions. However, they fall short in capturing the semantic correlation, as they learn high correlations for many other categories besides target category 674 (3rd row). The failure of existing methods to learn semantic-temporal correlation motivates us to develop a more effective user interest model. For a more comprehensive understanding of how we measure the learned correlation of each model, please refer to Sec. 4.3.
@@ 3. Temporal Interest Network
Figure 2. Left: the whole architecture of TIN; Right: the architecture of the Temporal Interest Module, which introduces target-aware temporal encoding, target-aware attention and representation to capture quadruple semantic-temporal correlations between behaviors and the target.
@@ 3.1. Temporal Interest Module
(2) $\bm{u}_{\text{TIN}}=\sum_{i\in\mathcal{H}}\alpha(\tilde{\bm{e}}_{i},\tilde{\bm{v}}_{t})\cdot(\tilde{\bm{e}}_{i}\odot\tilde{\bm{v}}_{t})$
@@ 3.1.2. Target-aware Attention and Target-aware Representation
Building upon target attention methods (Zhou et al., 2018, 2019; Feng et al., 2019), we aim to capture the importance of each behavior $X_{i}$ through an attention function. In this regard, we adopt the Scaled Dot-Product (Vaswani et al., 2017; Zheng et al., 2022) as our activation function: $\alpha(\tilde{\bm{e}}_{i},\tilde{\bm{v}}_{t})=\sigma(\frac{\langle\tilde{\bm{v}}_{t},\tilde{\bm{e}}_{i}\rangle}{\sqrt{d}})$, where $\sigma(z)$ denotes the softmax function on all behaviors of a given user.
@@ 3.3. Connections to Transformer
They differ from each other regarding 1. Scaled Dot-Product Attention is applied between each pair of tokens within the sequence, while TIM is only applied between the target and each item in the sequence. That is, Scaled Dot-Product Attention is a Self-Attention upon a sequence, while TIM is a Target-Attention upon a sequence and a target; 2. The representation in Scaled Dot-Product Attention is 1st-order only, that is, $V$, while it’s 2nd-order in TIM, that is, $(\tilde{\bm{e}}_{i}\odot\tilde{\bm{v}}_{t})$.
@@ 4. Experiments
RQ1: How does our proposed Temporal Interest Network (TIN) performance compare to state-of-the-art methods?
@@
RQ4: How do various temporal encoding methods perform compared to each other?
@@ 4.1. RQ1: Overall Performance Evaluation
The performance evaluation results are presented in Table 1. Avg Pooling & Concat exhibits the poorest performance on both datasets. This can be attributed to its lack of all three essential components, rendering it incapable of capturing either temporal correlation or semantic correlation (feature interaction). In contrast, Avg Pooling & Product demonstrates a significant performance improvement compared to Sum Pooling & Concat, with gains of 9.2e-3 and 5.3e-3 on each dataset, respectively. This improvement can be attributed to the inclusion of target-aware representation through the product operation, as discussed in (Rendle et al., 2020). While DIN’ performs worse than Avg Pooling & Product on three metrics, except for GAUC on the Alibaba dataset, this can be attributed to the limited expressiveness of Temporal Attention (TA) compared to TR. DIN outperforms DIN’ due to its incorporation of target-aware representation.
@@ 4.3. RQ3: Measurement of Learned Correlation
Figure 3. Learned semantic-temporal correlation of TIN and its three ablated variants on the Amazon dataset. The numbers in parentheses represent the Pearson correlation coefficient between the ground truth (as shown in Fig. 1(a)) and the learned correlation on behaviors with the same category as the target (3rd row).
@@ 4.4. RQ4: Temporal Encoding Methods
We partitioned the Amazon dataset based on the sequence length to evaluate these two temporal encoding methods and conducted a comparative analysis. As shown in Table 3, TTE-P consistently outperforms COE across all sequence length ranges, with the performance gap widening for longer sequences. Unfortunately, we are unable to perform a similar comparison on the Alibaba dataset since most sequences are padded to a fixed length of 50.
Next, we present the comparison results of TTE-P, TTE-T, and Timestamp Encoding on both datasets in Table 4. Specifically, TTE-T split the time interval into 10 buckets with an equal frequency binning function. Upon replacing TTE-P with TTE-T, we observe an increase in the GAUC by 8.6e-3 on the Amazon dataset and a slight drop of 0.8e-3 on the Alibaba dataset. This discrepancy can be attributed to the distinct time ranges of behaviors in the two datasets. The Amazon dataset encompasses review data from 1996 to 2014, whereas the Alibaba dataset comprises user behaviors recorded over a continuous period of 8 days. The larger time range of the Amazon dataset renders TTE-T more informative than TTE-P. Conversely, on the Alibaba dataset with a relatively short time range, TTE-P is more effective in capturing temporal patterns.
@@ 4.6. Online A/B Testing
We apply TIN to the user’s 2-year clicked ads category sequence feature in Tencent’s WeChat Moments and use both TTE-P and TTE-T to capture the temporal correlation. The backbone model employs a Heterogeneous Experts with Multi-Embedding architecture (Guo et al., 2023; Su et al., 2023b; Pan et al., 2024). Specifically, we learn multiple different feature interaction experts, e.g., GwPFM (Pan et al., 2024) (a variant of FFM (Juan et al., 2017) and FwFM (Pan et al., 2018)), IPNN (Qu et al., 2016), DCN V2 (Wang et al., 2021), or FlatDNN. There are hundreds of features, mostly user behavior features, ads, and context side features. Multiple embedding tables are learned for all features, each corresponding to one or several experts. TIN functions as an additional expert alongside the existing three experts, sharing embeddings with GwPFM and FlatDNN.
We conduct online A/B testing from September 2023 to October 2023. In the baseline, the user interest is modeled by feeding both user behavior features and target ads features to GwPFM or IPNN. We previously attempted to use DIN, but it did not surpass the baseline’s performance. This is possible because explicit interaction models such as GwPFM or IPNN can already learn the semantic correlation between behaviors and the target. During the two-week 20% A/B testing, TIN demonstrated promising results, achieving a 1.65% increase in cost and a 1.93% increase in GMV (Gross Merchandise Value) compared to the baseline. These improvements were statistically significant according to t-tests. Based on our estimation, these performance gains could potentially increase revenue by hundreds of millions of dollars per year. TIN has been successfully deployed on production in several scenarios.
@@ 6. Conclusion
Acknowledgements.
We gratefully acknowledge the contributions of the following: Xueming Qiu, Lei Mu, Xian Hu, Yaqian Zhang, Ming Yue, Wenbo Liu, Xiaobo Li, Zeen Xu, Jiayu Sun, Xun Liu, Yi Li, Ximei Wang, Junwen Cheng, Yan Tan, Yuxiong Li, Zhaohua Li, Kuo Zhang and Yufei Zheng. This work was supported by the National Natural Science Foundation of China [U23A20309, 62272302, 62172276, 62372296], Shanghai Municipal Science and Technology Major Project [2021SHZDZX-0102], and the Tencent Rhinoceros Project. Haolin Zhou, Xinyi Zhou, Xiaofeng Gao and Guihai Chen are in the MoE Key Lab of Artificial Intelligence, Department of Computer Science and Engineering, Shanghai Jiao Tong University, Shanghai, China. Xiaofeng Gao is the corresponding author.
@@ A.2. Expressiveness Study of Target-aware Attention and Target-aware Representation
Analysis of semantic-temporal correlation on more categories